OpenStudy (anonymous):

How do I show that f(y) = (k(theta^k))/y^(k+1), y >= theta is a valid probability density function?

OpenStudy (unklerhaukus):

is its sum of all probabilities equal to unity?

OpenStudy (anonymous):

if for all y, \[\int\limits f(y)dy=1\] then it will be a probability density function

OpenStudy (anonymous):

so if \[\int\limits_{-\infty}^{\infty}f(y)dy=\int\limits_{-\infty}^{\infty}k \theta^k/y^{k+1}dy=1\] then it will be a probability density function

OpenStudy (unklerhaukus):

find k such a that anonymoustwo44's equation's are true